What is the answer to 7 1/2 + 6 5/8 and how did you come up with the answer
Answer by ewatrrr(24785)

 
Hi
7 1/2 + 6 5/8 = 15/2 + 53/8 = (60+53)/8 = 113/8 = 14 1/8